### v1

# On Subjective Uncertainty Quantification and Calibration in Natural Language Generation

Abstract
Applications of large language models often involve the generation of free-form responses, in which case uncertainty quantification becomes challenging. This is due to the need to identify task-specific uncertainties (e.g., about the semantics) which appears difficult to define in general cases. This work addresses these challenges from a perspective of Bayesian decision theory, starting from the assumption that our utility is characterized by a similarity measure that compares a generated response with a hypothetical true response. We discuss how this assumption enables principled quantification of the model’s subjective uncertainty and its calibration. We further derive a measure for epistemic uncertainty, based on a missing data perspective and its characterization as an excess risk. The proposed measures can be applied to black-box language models. We demonstrate the proposed methods on question answering and machine translation tasks, where they extract broadly meaningful uncertainty estimates from GPT and Gemini models and quantify their calibration.
1 Introduction
We are interested in uncertainty quantification for language models (LMs) in free-form natural language generation (NLG): given instruction $I$, the model generates a response $y^{\prime}$ based on its predictive distribution $p_{M}(y\mid I)$, where $y^{\prime}$ could be any natural language passage that fits the instruction. An example of this is question answering (QA): given a question from the user, the model may provide a brief answer, but it may also follow with supporting facts and explanations, which can vary in form and detail. The user can be satisfied by a wide variety of responses, irrespective of their style or (to some extent) the choice of supporting facts included.
Free-form NLG poses significant challenges to uncertainty quantification: some aspects of generation are irrelevant to the task’s purpose and best excluded from uncertainty quantification, but it often appears that we are unable to characterize them precisely. If left unaddressed, however, the model’s variation in the irrelevant aspects may dominate in standard uncertainty measures such as token-level entropy (Kuhn et al.,, 2023), making them uninformative about the model’s actual performance on the task.
Starting from Kuhn et al., (2023), a recent line of work (Kuhn et al.,, 2023; Lin et al.,, 2024; Zhang et al.,, 2023; Aichberger et al.,, 2024) studied this issue and proposed measuring the “semantic uncertainty” of generation; “semantics” is defined as the equivalence class of textual responses that logically entail one another. Empirical improvements in downstream tasks evidenced their contributions and highlighted the importance of task-specific uncertainty quantification, but important conceptual and practical issues remain. From a practical perspective, semantic equivalence is estimated using machine learning models, resulting in imprecise estimates that do not necessarily define an equivalence relation. The imprecision necessitates the introduction of heuristics to post-process the estimates or to aggregate them through other means. Conceptually, the notion of semantic equivalence does not always provide a valid or complete characterization of relevance for certain tasks. Style transfer (Jin et al.,, 2022) tasks provide an example of the first kind. For the second scenario, consider a QA task where the model is fully certain about the answer to a question, as well as a large set of supporting facts, each of which provides complete and independent justification for the answer; yet there is still the “uncertainty”, or variation across generations, about which facts are included in a response. Then very few of the model generations may logically entail one another, and there is a very high level of “semantic uncertainty” just as if there was complete uncertainty about the correct answer.
This paper is about the observation that the above challenges can be resolved in a more general setup from a perspective of Bayesian decision theory (Savage,, 1954). We first assume that we can compute a similarity-type measure $S(y^{\prime},y;I)$ that characterizes our utility function when the model generates $y^{\prime}$ in respond to $I$ and $y$ is a (true or hypothetical) correct response. This generalizes previous works based on semantic equivalence, which corresponds to defining $S$ using entailment. It can also be applied to traditional structured prediction tasks where a choice of $S$ is readily available (e.g., the chrF score for machine translation, Popović,, 2015), as well as tasks where evaluation can be adequately implemented using frontier language models that computes $S$ based on few-shot demonstrations or detailed instructions.
To quantify the subjective uncertainty defined by the LM, we assume the generation $y$ is always chosen to maximize the expected utility under the LM’s predictive distribution $p_{M}$2.1). This is a common assumption in NLG (Bickel and Doksum,, 1997; Bertsch et al.,, 2023). It can be supported by a common belief that high-capacity LMs may approximate Bayesian inference (e.g., Xie et al.,, 2021; Akyürek et al.,, 2022; Hahn and Goyal,, 2023; Ye et al.,, 2024), and is also more broadly applicable whenever we view $p_{M}$ as our best available model for $y\mid I$. Subjective uncertainty is then naturally characterized by the Bayes risk, or equivalently the maximum achievable expected utility, where the action space is defined by the candidate generations available. This simple observation allows us to understand that previous methods for “semantic uncertainty” can be adapted to a broader range of scenarios, regardless of whether semantic equivalence is relevant. It also provides a unique, principled approach for aggregating similarity measures among generations (§2.2).
For subjective uncertainty measures to be useful, it will be ideal if the LM is calibrated. It has been unclear in previous works how calibration can be evaluated in free-form NLG tasks, given the distinction between relevant and irrelevant differences for a specific task and our apparent inability to formally characterise it. The decision-theoretic view provides a natural answer to this question: an LM is deemed as to possess a calibrated notion of uncertainty if the expected subjective utility of its actions matches the expectation of the actually incurred utility, defined w.r.t. responses from the true data distribution (§2.3). This observation allows us to quantify the calibration of LMs through reliability diagrams (Murphy and Winkler,, 1977) and a generalized version of expected calibration error (Naeini et al.,, 2015). In the discussion we also note a generally applicable definition for “task-specific semantics”: the equivalence class defined by the action-utility map, known as Savage acts.
Bayesian modelling allows for the decomposition of predictive uncertainty into epistemic uncertainty and aleatoric uncertainty (Der Kiureghian and Ditlevsen,, 2009). The decomposition allows us to better understand and diagnose the model’s uncertainty (e.g., only the epistemic uncertainty should decrease as the training set grows); it can also guide data acquisition (Kendall and Gal,, 2017). The quantification of epistemic uncertainty in general NLG applications appears challenging, due to the black-box nature of LMs and the free-form generations involved. §2.4 demonstrates that the decision-theoretic view enables us to quantify epistemic uncertainty for in-context learning (ICL, Brown et al.,, 2020) applications; key additional ingredients include a missing data perspective to Bayesian modelling (Fong et al.,, 2024) and a connection between epistemic uncertainty and a notion of excess risk (Xu and Raginsky,, 2022). The resulted measure is connected to and generalizes various previous works on epistemic uncertainty quantification for black-box models, which provides independent justification for its use.
The rest of this paper is structured as follows: §2 presents the proposed methodology and discusses connections to previous work; §3 provides experimental illustrations; and §4 provides concluding remarks.
2 Quantifying Subjective Uncertainty and Calibration
2.1 A Utilitarian Setup
The basic setup of this paper is as follows:
  • we have a NLG task: given prompt $I\in\mathcal{I}$, generate a response $y^{\prime}\in\mathcal{Y}$ where $\mathcal{Y}$ denotes the space of natural language responses;
  • our utility can be measured through the similarity measure $S(y^{\prime},y;I)\in\mathbb{R}$, which can be cheaply evaluated;
  • we have access to an LM $p_{M}$, and, in the absence of further evidence, consider $p_{M}$ sufficiently trustworthy so that the ideal generation (i.e., action) $y^{\prime}$ given $I$ should maximize the expected utility,
    $y^{\prime}:=\operatorname*{arg\,max}_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}S(y^{\prime},y;I),$ (1)
    where $\mathcal{Y}^{\prime}_{I}\subset\mathcal{Y}$ denotes the space of candidate generations.
In the above, $p_{M}$ represents a subjective belief about the true data distribution. $\mathcal{Y}_{I}^{\prime}\subset\mathcal{Y}$ can be determined based on computational constraints, e.g., as a separate set of samples from $p_{M}(\cdot\mid I)$.111 When $\mathcal{Y}_{I}^{\prime}$ is stochastic, expected utility/risk quantities such as (1) should also be averaged over the randomness in $\mathcal{Y}_{I}^{\prime}$. For brevity, we omit such (outmost) expectations throughout the paper. The utility $S$ determines a risk function $r(y^{\prime},y;I)=-S(y^{\prime},y;I)$, which can be scaled and shifted as desired.
As discussed in §1, this setup is very general. This decision-theoretic perspective is taken in the framework of minimum Bayes risk (MBR) decoding (Bickel and Doksum,, 1997), which advocates for generating $y$ based precisely on (1). MBR recovers many modern methods for natural language generation (Bertsch et al.,, 2023), and these methods can be viewed as having adopted the same view implicitly. Note that adopting the decision-theoretic setup does not compel us to assume that $p_{M}$ is a strictly Bayesian model (e.g., in the sense of Xie et al.,, 2021): although that will provide a good justification, it is also possible that a non-Bayesian LM may better approximate our belief than any available Bayesian model does.
2.2 Task-Specific Measure of Subjective Uncertainty
It is well known that Bayesian uncertainty measures are connected to minimum achievable risks of the form
$R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I}):=\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim p_{M}(\cdot\mid I)}r(y^{\prime},y;I).$ (2)
For example, if $\mathcal{Y}=\mathcal{Y}^{\prime}_{I}\subset\mathbb{R}$ and $r(y,y^{\prime};I)=(y-y^{\prime})^{2}$, $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ will be equivalent to the variance of the predictive distribution; if with an abuse of notation we redefine $y^{\prime}$ as a probability density function and $r(y^{\prime},y;I)\leftarrow\log y^{\prime}(y)$, $R_{B,\mathcal{Y}^{\prime}_{I}}(I)=H[p_{M}(\cdot\mid I)]$ will be the entropy. When the action space $\mathcal{Y}^{\prime}_{I}\subsetneq\mathcal{Y}$ is a strict subset, (2) will be connected to measures of usable information as discussed in Xu et al., (2019).
It is thus natural to also adopt (2) as a measure of subjective uncertainty for NLG tasks. To implement (2), we adopt Monte Carlo approximation for the expectation.
A common generation strategy is to follow the Gibbs predictor, which corresponds to (1) with $\mathcal{Y}^{\prime}_{I}:=\{y_{I}\}$ and $y_{I}\sim p_{M}(\cdot\mid I)$ being a single sample. In such cases, the expected uncertainty $\mathbb{E}_{\mathcal{Y}^{\prime}_{I}}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ is connected to the degree statistic in Lin et al., (2024, Eq. 8); the difference is that we use a similarity measure that is assumed to define our utility. While Lin et al., (2024) explored 9 different uncertainty measures for the Gibbs predictor, the decision-theoretic perspective leads to one unique method without any underspecified design choices, justifies its use beyond the scope of semantic equivalence, and highlights that different choices of predictors ($\mathcal{Y}_{I}^{\prime}$) should be matched with different uncertainty measures. However, being a measure of subjective uncertainty, it is not guaranteed to always improve on downstream tasks compared with heuristic alternatives, especially if the LM is less calibrated.
2.3 Evaluation of Task-Specific Calibration
Subjective uncertainty measures are most useful when the LM is calibrated. As discussed in Huang et al., (2024), many previous works studying uncertainty on NLG tasks did not distinguish between calibration and predictive performance in their evaluation, which leads to a relative lack of understanding about the calibration of LMs in such scenarios.
A typical definition of calibration is as follows (see e.g., Johnson et al.,, 2024): suppose the spaces $\mathcal{Y},\mathcal{I}$ are discrete, and denote the true data distribution as $p_{0}$; a model $p_{M}$ is considered calibrated w.r.t. some grouping function $G(I):\mathcal{I}\to\mathcal{G}$ if we have
$\mathbb{E}_{p_{0}}(p_{M}(y=y^{\prime}\mid I)\mid G(I)=g)=\mathbb{E}_{p_{0}}(p_{0}(y=y^{\prime}\mid I)\mid G(I)=g)~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}.$ (3)
The grouping function $G$ may be defined based on the model’s prediction, or it may incorporate additional information such as input demographics. In the NLG setting, however, the criterion (3) can be trivially violated if, e.g., the LM always generates a longer response compared with $p_{0}$.
In light of the assumptions in §2.1, it is natural to relax (3) as follows:
$\mathbb{E}(p_{M}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$ $=\mathbb{E}(p_{0}(\{y:S(y^{\prime},y;\mathcal{I})=s\}\mid I)\mid G(I)=g)$
$\hskip 130.0002pt~{}~{}\forall g\in\mathcal{G},y^{\prime}\in\mathcal{Y}^{\prime}_{I},s\in\mathbb{R}.$ (3’)
(3) is equivalent to (3) if $\mathcal{Y}^{\prime}_{I}=\mathcal{Y}$ and $S(y^{\prime},y;\mathcal{I})=\mathbf{1}\{y=y^{\prime}\}$, or more generally if for all $I\in\mathcal{I}$ the map from $y$ to the function
$s_{y,I}:\mathcal{Y}_{I}^{\prime}\to\mathbb{R},~{}y^{\prime}\mapsto S(y^{\prime},y;I)$ (4)
is injective (e.g., if $S$ is defined by a metric on $\mathcal{Y}$). (3) becomes weaker if $S(y^{\prime},y;\mathcal{I})$ only depends on $y$ and $y^{\prime}$ through a non-invertible function, e.g., a map from a natural language passage to its semantics; in such cases, it is clear that (3) remains a sufficiently strong criterion for task-specific calibration. More generally, we can always view the function $s_{y,I}(y^{\prime})$ as a representation of “task-specific semantics”.222 For the goals of providing the best generation from $\mathcal{Y}^{\prime}_{I}$ and quantifying its risk, it suffices to define the function $s_{y,I}$ on $\mathcal{Y}^{\prime}_{I}$. For purposes beyond the scope of the generation task, it may be helpful to extend its domain to a larger subset of $\mathcal{Y}$. In fact, the replacement of $y^{\prime}\in\mathcal{Y}^{\prime}_{I}$ with the more compact $s_{y^{\prime},I}(\cdot)$ has been advocated by Savage, (1954), and $s_{y,I}$ is referred to as a Savage act (see e.g., Marinacci,, 2015, §2.4).
A main reason we are interested in criteria such as (3) or (3) is that they guarantee the calibration of uncertainty measures. It follows from (3) that $\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(y^{\prime},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(y^{\prime},y;I)$ for all $y^{\prime}$, and thus
$\mathbb{E}_{p_{0}(I\mid G(I)=g)}R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{M}(y\mid I)}r(\hat{y}_{I},y;I)=\mathbb{E}_{p_{0}(I\mid G(I)=g)p_{0}(y\mid I)}r(\hat{y}_{I},y;I),$ (5)
where $\hat{y}_{I}$ denotes the MBR generation (1). Plugging in $G(I)=R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$, we find that LMs are calibrated in the sense of (3) only if
$\text{gECE}(p_{M}):=\mathbb{E}_{s}|f_{M}(s)-s|=0,~{}~{}~{}\text{where }f_{M}(s):=\mathbb{E}_{I,y,\hat{y}_{I}}(r(\hat{y}_{I},y;I)\mid R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})=s)~{}\forall s\in\mathbb{R}.$ (6)
(6) provides a natural generalization of the expected calibration error (ECE, Naeini et al.,, 2015) in classification, which it recovers with $r(y^{\prime},y;I)\leftarrow\mathbf{1}\{y^{\prime}\neq y\}$. Similar to the classification case, (6) can be estimated through histogram binning. The plot for the function $f_{M}$ is called a reliability diagram (Murphy and Winkler,, 1977) and provides information on overconfidence ($f_{M}(s)<s$) or underconfidence. Importantly, these methods remain applicable in NLG as we have shown. Moreover, they will automatically focus on the task-specific aspects of uncertainty, since the function $f_{M}$ only depends on model generations through the utility $S$.
Remark 2.1.
We discussed the evaluation of LM calibration based on the uncertainty measure (2). To our knowledge, only Huang et al., (2024) studied a similar problem in the general NLG setting, namely the rank-calibration of general uncertainty measures. As discussed in Huang et al., (2024, §4.2), rank-calibration is still a different property; in particular, (2) will be rank-calibrated whenever $f_{M}$ is monotone, even though it can be arbitrarily overconfident or underconfident and thus miscalibrated.
2.4 Representing and Eliciting Epistemic Uncertainty
We now turn to the problem of epistemic uncertainty quantification in in-context learning, where the prompt $I=\langle x_{1},y_{1},\ldots,x_{n},y_{n},x_{*}\rangle$ is formed by $n$ demonstrations $(x_{i},y_{i})$ followed by a test input $x_{*}$. We assume that $r(y,y^{\prime};I)=r(y,y^{\prime};x_{*})$ only depends on $I$ through $x_{*}$, which is commonly satisfied in applications. We propose the following measure of epistemic uncertainty inspired by Fong et al., (2024); Xu and Raginsky, (2022):
$\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*})$
$\hskip 40.00006pt-\mathbb{E}_{p_{x}(x_{n+1:N})}\mathbb{E}_{p_{M}(y_{n+1:N}\mid y_{1:n},x_{1:N})}\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(y_{*}\mid y_{1:N},x_{1:N},x_{*})}r(y^{\prime},y_{*};x_{*}).$ (7)
In the above, $p_{x}$ denotes a user-specified distribution of additional ICL inputs, $y_{n+1:N}$ are model generations, $N>n$ is a hyperparameter, and we assume the LM has good performance for ICL up to a sample size of $N$. Eq. (7) can be estimated through autoregressive sampling. As long as we use the same set of samples for its both terms, the resulted estimate will always be non-negative, due to the convexity of the functional $\rho\mapsto\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{\rho(y_{*})}r(y^{\prime},y_{*};x_{*})$.
To understand (7), first consider a simplified scenario as follows: suppose the LM $p_{M}$ is equivalent to a Bayesian model that assumes $\{(x_{i},y_{i})\}$ are i.i.d. conditional on a latent variable $\theta$, $p_{x}$ defines i.i.d. $x_{n+1:N}$, and $\{x_{n+i},y_{n+i}\}$ are sufficiently informative so that the likelihood function $p(y=\cdot\mid x=x_{*},\theta)$ determined by $\theta\mid x_{1:N},y_{1:N}$ is identifiable given infinite samples. Then under mild technical conditions, as $N\to\infty$ (7) will become equivalent to
$\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*})-\mathbb{E}_{p_{M}(\theta\mid x_{1:n},y_{1:n})}\min_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{p(y\mid\theta,x=x_{*})}r(y^{\prime},y;x_{*}).$ (7’)
The equivalence is a consequence of Doob, (1949)’s theorem and can be proved with the same idea as Fong et al., (2024). For completeness, we provide a proof in Appendix A. (7) is a subjective notion of excess risk. In other words, it is the amount of reducible risk should we have full knowledge about the distribution of $y\mid x=x_{*}$, which can be determined given either the latent $\theta$, or an infinite amount of missing data $\{x_{>n},y_{>n}\}$. As noted in Xu and Raginsky, (2022), this is precisely the intuition behind Bayesian epistemic uncertainty, and we recover standard epistemic uncertainty measures (e.g., posterior variance for a regression mean, or mutual information) if we return to simple scenarios where $r$ is the square or log loss.
Within the Bayesian framework, (7) with a finite choice of $N$ will provide a lower bound for (7): by convexity the second term may become larger. The intuition is that knowledge of the samples $\{x_{n+1:N},y_{n+1:N}\}$ does not always allow for the full reduction of risk. It is possible to view the lower bound as an indication of the “true epistemic uncertainty” (7), and it can be tightened if we choose $\{x_{n+i}\}$ to be more similar to $x_{*}$. The bound will also be tight regardless of the choice of $N$ if the uncertainty is “fully aleatoric” or “fully epistemic”, i.e., if the risk for $y\mid x=x_{*}$ cannot be reduced by any number of additional observations $(x_{n+1:N},y_{n+1:N})$ or can be fully reduced using a single $(x_{n+1},y_{n+1})$. Alternatively, we can observe that we do not always need the tightest bound for (7): if all we can do is to collect $N-n$ real samples, the remaining proportion of the risk will be effectively irreducible.
As discussed in §1, the Bayesian view above may be justified through a common belief that high-capacity LMs may approximate Bayesian inference. It is also supported by the results of Wen et al., (2022, §2.6, §4), which imply that in a multi-task setting, any $p_{M}$ with a strong average-case performance for prediction will yield a similar value for (7) as the true Bayesian model. Falck et al., (2024) investigated this belief empirically on synthetic low-dimensional datasets, and found it to be valid for smaller choices of $N$, although it breaks down as $N\to\infty$ becomes larger.333 Falck et al., (2024) also studied epistemic uncertainty quantification for ICL but restricted to datasets where the method of Fong et al., (2024) is applicable. Their method assumes $\{x_{n+1:N},$ $y_{n+1:N}\}$ can be modelled by a low-dimensional statistical model, so that a posterior for the model parameter can be constructed and visualized. This is unrealistic for free-form natural language data. Taken together, these results suggest that the Bayesian perspective could be relevant when we restrict $N$ to a scope where ICL is sample efficient.
Independent to the Bayesian perspective, Eq. (7) can also be understood through its connection to recent works. Collier et al., (2022) studied a broadly similar problem where a proportion of the combined uncertainty can be reduced by conditioning on privilege information; this leads to a similar measure for reducible uncertainty as (7), if we view $(x_{n+1:N},y_{n+1:N})$ as the latent “privilege information”. Johnson et al., (2024); Ahdritz et al., (2024) proposed methods that measure the correlation between consecutive predictions $(y_{n+1},y_{*})$ given the same input $x_{n+1}=x_{*}$; as argued in Johnson et al., (2024), any learning system $p_{M}$ with complete confidence for the true data distribution should lead to conditionally i.i.d. $(y_{*},y_{n+1})$ given $I$, and any $p_{M}$ with a “fully epistemic” uncertainty should produce $y_{*}=y_{n+1}$. It is clear that (7) matches these behaviours at the extremes: it equals 0 in the former case, and correctly indicate a full reduction of risk in the latter case (i.e., it equals $R_{B}(I;p_{M},\mathcal{Y}^{\prime}_{I})$ assuming $r(y,y;I)\equiv 0$).The Bayesian perspective provides additional insights to the works of Johnson et al., (2024); Ahdritz et al., (2024). Eq. (7) generalizes their methodology by allowing for a wider range of choices for $\{x_{n+1:N}\}$, and the above discussion provides guidance on their choices.
Finally, we note that while (7) is generally applicable to MBR decoding, with a trivial choice of $\mathcal{Y}_{I}^{\prime}$ it may take a trivial value: if we use the Gibbs predictor which always samples from $p_{M}(\cdot\mid x_{1:n},y_{1:n})$, we will find (7) to be zero; this is to be expected as the predictor does not make any attempt to utilize the additional information. In our experiments we choose $\mathcal{Y}_{I}^{\prime}$ to contain multiple samples, one from each sample path for $y_{n+1:N}$. As we will see, this is sufficient to provide meaningful estimates of epistemic uncertainty.
3 Experiments
3.1 Free-Form Question Answering
We first illustrate the proposed method on two QA tasks chosen by Lin et al., (2024): CoQA (Reddy et al.,, 2019) and NQOpen (Lee et al.,, 2019). CoQA contains reading comprehension questions where the answer can generally be deduced from the prompt, whereas NQOpen is an open-domain task and requires information not present in the prompt. Kalai and Vempala, (2024) proved in simplified scenarios that instruction-tuned LMs may have to be uncalibrated on the latter type of task; we investigate their theoretical predictions empirically.
Figure 1: Reliability diagrams for GPT-3.5 on QA tasks. Error bar denotes 95% confidence interval for the average observed utility.
We follow the setup in Lin et al., (2024) and adopt the GPT-3.5 model (OepnAI,, 2024), both for generation and to define the automated evaluation criterion; the latter determines our utility function and in turn the uncertainty measure (2).444 In this aspect our method differs from Lin et al., (2024), where uncertainty measures are defined with estimated entailment relations or lexical metrics. The evaluation criterion is defined using ICL, by instructing the LM to rate the consistency between two responses based on few-shot demonstrations (Lin et al.,, 2024, App. B.3). We reuse the model samples released by Lin et al., (2024), which contains 20 sampled responses for each query. Following their work we focus on the Gibbs predictor.
The results are shown in Fig. 1. As we can see, the LM’s subjective uncertainty generally correlates well with the average observed utility. It is slightly overconfident on the reading comprehension task, but much more so on the open-domain task. Such a difference appears consistent with the theoretical analysis of Kalai and Vempala, (2024).
3.2 In-Context Machine Translation
We now consider machine translation with ICL, where utility is defined through the widely adopted chrF score (Popović,, 2015). chrF measures both semantic and syntactic differences; the latter can be relevant in translation due to the need to detect unnatural or invalid syntactic structures. The inherent syntactic differences in the possible generations for $y\mid x$ constitute a source of irreducible uncertainty, making machine translation a non-trivial scenario to illustrate our uncertainty decomposition method.
We follow the setup of Agarwal et al., (2024) who demonstrated that ICL can lead to steady improvements for machine translation. We adopt the FLORES+ dataset (NLLB Team et al.,, 2022), in particular the translation tasks from English to the following languages: French, Tamil, Yue Chinese (Cantonese), and Tigrinya. We use the following LMs: GPT-3.5, Gemini-1.5-Flash, and Gemini-1.5-Pro (Gemini Team,, 2024), and vary the number of ICL shots $n\in\{4,16\}$. For each task, we subsample 128 queries $(x_{*},y_{*})$ from the datasets. For each query, we draw 35 samples from the LM to approximate its predictive distribution, and another 5 samples to define the action space $\mathcal{Y}^{\prime}_{I}$. For the quantification of epistemic uncertainty, we set $N=n+8$, and sample $x_{n+1:N}$ by instructing an LM to rewrite the test query $x_{*}$; see Appendix B for details.
$n$ gemini-1.5-flash gemini-1.5-pro gpt-3.5
$4$ 0.276 / 0.356 0.332 / 0.431 0.185 / 0.329
$16$ 0.246 / 0.371 0.304 / 0.426 0.168 / 0.351
Table 1: Machine translation (English $\to$ Yue Chinese): gECE (Eq. (6), $\downarrow$) and chrF ($\uparrow$) for varying choices of $n$.
We first evaluate the calibration of the combined subjective uncertainty on the Yue Chinese task. As shown in Table 1, the LMs are generally poorly calibrated, although calibration improves as the sample size $n$ increases. Calibration plots are presented in Figure 3 in Appendix, which shows that all LMs are overconfident. It is interesting to note an anticorrelation between the task performance and calibration among the three LMs evaluated. We note that Yue Chinese may constitute a low-resource translation task (Haddow et al.,, 2022) for which the LMs’ pretraining corpus is likely less informative. As is also shown in Figure 3, calibration is better on the French task for which more corpus is available.
Figure 2: Machine translation (English $\to$ Yue Chinese): reduction of epistemic uncertainty (EU) vs total uncertainty (TU) as the ICL sample size increases from $4$ to $16$.
We now investigate whether the LMs possess a coherent notion of epistemic and aleatoric uncertainty, in particular whether epistemic uncertainty is the only reducible component as sample size increases. For this purpose we plot the reduction of total uncertainty (TU, (2)) against that of epistemic uncertainty (EU, (7)) on each data point. Figure 2 shows that the two quantities have a good match. Deviation from perfect equivalence is expected here, due to the variance in the Monte-Carlo estimation for EU and TU.
Language French Tamil Yue Tigrinya
Avg. EU 0.016 0.059 0.078 0.091
chrF 0.737 0.392 0.329 0.084
# speakers (M) 312 87 87 10
Table 2: Machine translation: average epistemic uncertainty on different languages, obtained from gpt-3.5 using $n=4$ ICL shots. We also provide the chrF score and the number of speakers (from Ethnologue,, 2024, in million) of each language as an indication of its linguistic resource available.
Finally, we evaluate if the epistemic uncertainty measure (7) is meaningful, by comparing its average value across different translation tasks: we expect sensible notions of epistemic uncertainty to be higher on tasks for low-resource languages, as they are likely less represented in the pretraining corpus. Table 2 confirms this intuition: the LM demonstrates higher epistemic uncertainty on lower-resource languages.
4 Conclusion
This work studies uncertainty quantification in free-form natural language generation. Through a Bayesian decision-theoretic perspective, we derived principled methods to quantify the model’s task-specific subjective uncertainty and to evaluate its calibration. The proposed methods are also connected to previous work, which provides additional justification for their adoption. While the discussions (in particular those in §2.2 and §2.3) are not necessarily novel outside the context of language modelling, within this domain the decision-theoretic perspective appears new and addresses important conceptual challenges in uncertainty quantification. Experiments demonstrate that the methods extract meaningful uncertainty estimates from black-box LMs and quantify their calibration.
Our discussion concerns the quantification of subjective uncertainty. In applications where i.i.d. calibration samples are available, it would be beneficial to recalibrate (Marx et al.,, 2022) the uncertainty measures, which can be achieved through (6), or to incorporate the uncertainty measures in conformal prediction (Quach et al.,, 2023). As the capabilities of LMs continue to evolve, it is also conceivable that future-generation models may find use in “one-of-its-kind” applications where post-calibration and conformal approaches are inapplicable. In such scenarios, it would be crucial to have principled measures for subjective uncertainty, as well as the tools to dissect and understand the sources of this uncertainty, e.g., through the proposed decomposition into epistemic and aleatoric uncertainty. It would also be interesting for future work to investigate the consistency between the LM’s verbalized uncertainty (Kadavath et al.,, 2022) and the subjective uncertainty defined by its predictive distributions.
Appendix A Proof for the Equivalence between (7) and (7)
Claim A.1.
Suppose $\mathcal{X},\mathcal{Y}$ are finite sets, and $\sup_{y^{\prime},y,x}|r(y^{\prime},y;x)|<\infty$. Suppose there exists a Bayesian model with prior $\pi$ over a discrete parameter $\theta\in\Theta$ and a (conditional) likelihood function $p(y\mid x,\theta)$, s.t. for all $(n,x_{1:n},y_{1:n},x_{*})$ we have
$p_{M}(y_{n+1}=\cdot\mid x_{n+1}=x_{*},x_{1:n},y_{1:n})=\int\pi(d\theta\mid x_{1:n},y_{1:n})p(y=\cdot\mid x=x_{*},\theta)$
where $\pi(d\theta\mid x_{1:n},y_{1:n})\propto\pi(d\theta)\prod_{i=1}^{n}p(y=y_{i}\mid x=x_{i},\theta)$ denotes the parameter posterior. Suppose (7) is defined using $x_{n+1:N}\overset{i.i.d.}{\sim}p_{x,1}$, and that for all $x_{*}\in\mathcal{X}$ and $\pi$-a.e. $\theta$ we have
$\lim_{n\to\infty}\mathbb{E}_{x_{1:n}\sim p_{x,1},y_{i}\sim p(y=\cdot\mid x=x_{i},\theta)}\|p_{M}(y_{n+1}=\cdot\mid x_{n+1}=x_{*},x_{1:n},y_{1:n})-p(y=\cdot\mid x=x_{*},\theta)\|_{\ell_{2}(\mathcal{Y})}^{2}=0,$ (8)
where $\|f\|_{\ell_{2}(\mathcal{Y})}:=\sqrt{\sum_{y\in\mathcal{Y}}f(y)^{2}}$ denotes the $\ell_{2}$ norm. Then for $p_{x,1}$-a.e. $(x_{1:n},x_{*})$ and $p_{M}$-a.e. $y_{1:n}$, the $n\to\infty$ limit of (7) is equivalent to (7).
In the above, (8) is the assumed identifiability condition for the likelihood function. We expect the numerous technical restrictions to be relaxable, but refrain from a more general proof for brevity. The main conditions that enable the equivalence are that (i) $p_{M}$ can be augmented to define an exchangeable (or c.i.d.) model for $(x,y)$, and (ii) the likelihood function is identifiable.
Proof.
We will show that each of the two terms in (7) and (7) are equivalent.
Let us introduce the following notations: define $\mathcal{Z}:=\mathcal{X}\times\mathcal{Y}$ and, for any $x_{i},y_{i}$, $z_{i}:=(x_{i},y_{i})\in\mathcal{Z}$. For all $n\geq 0$, define $\bar{p}_{M}(z_{n+1}=(x,y)\mid z_{1:n})=p_{x,1}(x)p_{M}(y_{n+1}=y\mid x_{n+1}=x,x_{1:n},y_{1:n}),\bar{p}(z=(x,y)\mid\theta)=p_{x,1}(x)p(y\mid\theta,x)$. Clearly, $\bar{p}_{M}$ is equivalent to a Bayesian model with prior $\pi$ and the factorized likelihood $\bar{p}$. Thus, $z_{n+1:N+1}\sim\bar{p}_{M}(\cdot\mid z_{1:n})$ are exchangeable, and we have, for all $n,z_{1:n},N>n,y^{\prime}\in\mathcal{Y}$ and $p_{x,1}$-almost every $x_{*}$,
$\mathbb{E}_{\bar{p}_{M}}(r(y^{\prime},y_{N+1};x_{*})\mid z_{1:n},x_{N+1}=x_{*})=\mathbb{E}_{\bar{p}_{M}}(r(y^{\prime},y_{n+1};x_{*})\mid z_{1:n},x_{n+1}=x_{*}).$
By definitions, the terms above equal the first term in (7) and (7), respectively, with their outmost infimum removed. Retaking infimum proves the equivalence of the first terms.
For the second term, observe that by the boundedness of $r$ and the discreteness of $\mathcal{Y}$, for all $x_{*}\in\mathcal{X}$ the function $R_{x_{*}}(\rho):=\inf_{y^{\prime}\in\mathcal{Y}^{\prime}_{I}}\mathbb{E}_{y\sim\rho}r(y^{\prime},y;x_{*})$ is Lipschitz continuous w.r.t. the $\ell_{2}(\mathcal{Y})$ norm. [^5] Note that any distribution $\rho$ over $\mathcal{Y}$ can be identified with a probability mass function in $\ell_{2}(\mathcal{Y})$ since $\mathcal{Y}$ is discrete. Thus, for $\pi$-a.e. $\theta$ we have
$\lim_{N\to\infty}\mathbb{E}_{z_{1:N-n}\overset{i.i.d.}{\sim}\bar{p}(z\mid\theta)}|R_{x_{*}}(p_{M}(y=\cdot\mid x=x_{*},z_{1:N-n}))-R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta))|=0$
following (8). It follows by the boundedness of $R_{x_{*}}(\cdot)$ and the dominated convergence theorem that
$\phantom{=}\lim_{N\to\infty}\mathbb{E}_{\theta\sim\pi,z_{1:N-n}\sim\bar{p}(z\mid\theta)}R_{x_{*}}(p_{M}(y=\cdot\mid x=x_{*},z_{1:N-n}))=\mathbb{E}_{\theta\sim\pi}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta)).$ (9)
The RHS above equals the expectation of the second term in (7) over $z_{1:n}\sim\bar{p}_{M}$. Denote the second term in (7) as $u_{N}(z_{1:n})$. By definition of $\bar{p}_{M}$ and its exchangeability, as well as the convexity of $R_{x_{*}}(\cdot)$, we have
$\mathbb{E}_{\bar{p}_{M}(z_{1:n})}u_{N}(z_{1:n})$ $=\mathbb{E}_{\bar{p}_{M}(z_{1:n})}\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}R_{x_{*}}(\bar{p}_{M}(y_{N+1}=\cdot\mid x_{N+1}=x_{*},z_{1:N}))$
$\leq\mathbb{E}_{\bar{p}_{M}(z_{n+1:N})}R_{x_{*}}(\mathbb{E}_{\bar{p}_{M}(z_{1:n}\mid z_{n+1:N})}\bar{p}_{M}(y_{N+1}=\cdot\mid x_{N+1}=x_{*},z_{1:N}))$
$=\mathbb{E}_{\bar{p}_{M}(z_{1:N-n})}R_{x_{*}}(\bar{p}_{M}(y_{N-n+1}=\cdot\mid x_{N-n+1}=x_{*},z_{1:N-n}))$
$=\mathbb{E}_{\theta\sim\pi,z_{1:N-n}\sim\bar{p}(z\mid\theta)}R_{x_{*}}(p_{M}(y=\cdot\mid x=x_{*},z_{1:N-n})).$
Combining with (9) and applying the dominated convergence theorem yield
$\mathbb{E}_{\bar{p}_{M}(z_{1:n})}\lim_{N\to\infty}u_{N}(z_{1:n})=\lim_{N\to\infty}\mathbb{E}_{\bar{p}_{M}(z_{1:n})}u_{N}(z_{1:n})\leq\mathbb{E}_{\bar{p}_{M}(z_{1:n})}\mathbb{E}_{\pi(\theta\mid z_{1:n})}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta)).$ (10)
On the other hand, we have
$u_{N}(z_{1:n})$ $=\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}R_{x_{*}}(\bar{p}_{M}(y=\cdot\mid x=x_{*},z_{1:N}))$
$=\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}R_{x_{*}}(\mathbb{E}_{\pi(\theta\mid z_{1:N})}p(y=\cdot\mid x=x_{*},\theta))$
$\geq\mathbb{E}_{\bar{p}_{M}(z_{n+1:N}\mid z_{1:n})}\mathbb{E}_{\pi(\theta\mid z_{1:N})}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta))$
$=\mathbb{E}_{\pi(\theta\mid z_{1:n})}R_{x_{*}}(p(y=\cdot\mid x=x_{*},\theta)),$
so the same holds for $\lim_{N\to\infty}u_{N}(z_{1:n})$. Comparing with (10) we find that equality must hold for $\bar{p}_{M}$-a.e. $z_{1:n}$. This proves the equivalence for the second terms in (7) and (7), and consequently the original claim. ∎
Appendix B Additional Experimental Details
In both experiments we use the following versions of LM APIs: gpt-3.5-turbo-0125, gemini-1.5-flash-001, gemini-1.5-pro-001.
The machine translation experiment is implemented as follows: for each test query $x_{*}$, we sample the ICL demonstrations $(x_{1:n},y_{1:n})$, a single set of $x_{n+1:N}$, and $5$ sets of $y_{n+1:N}\sim p_{M}(x_{1:N},y_{1:n})$; then for each path sample $y_{n+1:N}$ we draw $8$ samples from $p_{M}(y_{*}\mid x_{1:N},y_{1:N},x_{*})$. The action space is defined using $5\times 1$ samples from each path sample, and the expectations in (2) and (7) using the remaining $5\times 7$ samples. For each $x_{*}$, we reuse the same set of samples for $(x_{n+1:N},x_{1:n},y_{1:n})\mid x_{*}$ across all experiments. The completion inputs $x_{n+1:N}$ are generated using Gemini-1.5-Flash with the following prompt:
Please generate {K} sentences that convey a broadly similar meaning as the one provided, ensuring they are of similar length and use varied wording. Place a blank line after each sentence. Avoid adding any extra explanations.
{x_*}
In the above, $x_{*}$ denotes the test query. For simplicity, we report the chrF scores evaluated using the final samples $y_{*}\mid x_{1:N},y_{1:N},x_{*}$, which is similar to the reinforced ICL method (Agarwal et al.,, 2024).
Figure 3: Machine translation: reliability diagrams for a subset of experiments. Error bar indicates 95% confidence interval for the expected observed utility.